OpenStudy (anonymous):

Could someone please explain this? A 14-ft-long ramp rises at an angle of 22.2°. How long is the base of the ramp to the nearest foot? (1 point) 11 ft 13 ft 17 ft 22 ft

OpenStudy (stamp):



OpenStudy (anonymous):

Still don't get it.

Directrix (directrix):

cosine of (22.2 degrees) = b / 14 Solve for b. @Channi

Directrix (directrix):

cosine of (22.2 degrees) = ? @Channi

OpenStudy (anonymous):

0.92 apparently. Okay well that's wrong, I'll go ahead and give you a metal. If I'm too stupid see what I assume is obvious here, surely I deserve to get it wrong. Good day to you people.

Directrix (directrix):

cosine of (22.2 degrees) = b / 14 The choice of the cosine function to solve this problem comes from the diagram which @stamp drew above.

Directrix (directrix):

Enter "cosine of 22.2 degrees" in the Google search box, and the Google calculator will pop up and render 0.9258 (approx) as the result.

Directrix (directrix):

cosine of (22.2 degrees) = b / 14 .9258 = b/14 b = .9258* 14 b = 12.97 or 13 feet to the nearest foot @Channi

OpenStudy (anonymous):

Wow! That was a lot simpler than I thought it was. Was I seriously not substituting. Damn. Thank you :)

Directrix (directrix):

We were glad to help, @Channi .